Find the frequency of the 3rd harmonics of a closed pipe 400cm long if the velocity of sound in air is 350ms/1


1
Expert's answer
2021-07-28T09:34:04-0400

The harmonics in the closed pipe are given as follows (see https://en.wikipedia.org/wiki/Standing_wave#Standing_wave_in_a_pipe):


"f_n = \dfrac{mv}{2L}"

where "n = 1, 2, 3..." is a positive integer, "v = 350m\/s" is the speed of sound, and "L = 400cm = 4m" is the length of the pipe. Substituting "n = 3", obtain:


"f_3 = \dfrac{3\cdot 350m/s}{2\cdot 4m} \approx 131Hz"

Answer. 131 Hz.